OpenStudy (anonymous):

Need help on figuring this answer: ln(3) - 2ln(4+8) I got 3.9 and wrong

OpenStudy (lalaly):

do u knw the answer?

OpenStudy (anonymous):

no I don't that'st why I'm asking not sure what I did wrong

OpenStudy (lalaly):

i got ln3-2ln12 = ln3-ln12^2 = ln3-ln44 = ln(3/44) = -2.68

OpenStudy (anonymous):

From Mathematica: (Ln is spelled Log below.) Log[3] - 2*Log[4 + 8] = -3.8712

OpenStudy (anonymous):

try this \[\ln(3)-\ln(12)=\ln(\frac{1}{4})\]

OpenStudy (anonymous):

are you just looking for the number?

OpenStudy (lalaly):

you forgot the 2ln12 satellite